OpenStudy (anonymous):

Suppose that f(x)=sqrt(4+sqrt(3x)) Find f'(x) Using the Chain rule Please help I have been trying to figure this out all day. So far I have this (1/2(4+sqrt3)^-1/2) (1/2(sqrt3)^-1/2)

OpenStudy (campbell_st):

ok you need to rewrite you function in index notation to make it easier... \[y= ( 4 +(3x)^{\frac{1}{2}})^{\frac{1}{2}}\] so now let \[u = 4 + (3x)^{\frac{1}{2}}\] then \[y = u^{\frac{1}{2}}\] find \[\frac{dy}{du}\] also find \[\frac{du}{dx} \] then \[f'(x) = \frac{dy}{du} \times \frac{du}{dx}\]

OpenStudy (campbell_st):

so you solution looks correct so far \[\frac{1}{2}( 4 + \sqrt(3x))^{-\frac{1}{2}} \times \frac{1}{2}(3x)^{-\frac{1}{2}}\] now you need to be able to write it without index notation. the 1st part is \[\frac{1}{2} \times \frac{1}{\sqrt{4 + \sqrt{3x}}}\]

OpenStudy (anonymous):

So would the next one be 1/2*(1/sqrt3x)?

OpenStudy (campbell_st):

correct...

OpenStudy (campbell_st):

so you will have \[\frac{1}{2\sqrt{4 + \sqrt{3x}}} \times \frac{1}{2\sqrt{3x}}\] just simplify this

OpenStudy (anonymous):

This isn't be accepted as the correct answer.